OpenStudy (trojanpoem):

Derivatives question:

OpenStudy (trojanpoem):

\[y = (x+ \sqrt{x^2 +1 })^4\]

OpenStudy (trojanpoem):

Prove that : \[(x^2 + 1)y ^{''} + xy ^{'} = 16y\]

OpenStudy (michele_laino):

here we have to compute the first and second derivative of y(x)

OpenStudy (trojanpoem):

It got complicated much !

OpenStudy (michele_laino):

no, it is very simple, please note that: \[\Large y' = \frac{{4y}}{{\sqrt {{x^2} + 1} }}\]

OpenStudy (trojanpoem):

Oo, But how did you get rid of 4 power ?

OpenStudy (trojanpoem):

Hey ?

OpenStudy (michele_laino):

here are your steps: \[\Large \begin{gathered} y' = 4{\left\{ {x + {{\left( {{x^2} + 1} \right)}^{1/2}}} \right\}^3}\left\{ {1 + \frac{x}{{{{\left( {{x^2} + 1} \right)}^{1/2}}}}} \right\} = \hfill \\ \hfill \\ = 4{\left\{ {x + {{\left( {{x^2} + 1} \right)}^{1/2}}} \right\}^3}\frac{{x + {{\left( {{x^2} + 1} \right)}^{1/2}}}}{{{{\left( {{x^2} + 1} \right)}^{1/2}}}} = \hfill \\ \hfill \\ = 4\frac{{{{\left\{ {x + {{\left( {{x^2} + 1} \right)}^{1/2}}} \right\}}^4}}}{{{{\left( {{x^2} + 1} \right)}^{1/2}}}} = \frac{{4y}}{{\sqrt {{x^2} + 1} }} \hfill \\ \end{gathered} \]

OpenStudy (trojanpoem):

Wow ! I didn't think about the 2nd step. But now , do I get the dev directly or ^2 each ?

OpenStudy (michele_laino):

now, we can compute the second derivative, applying the quotient rule

OpenStudy (michele_laino):

\[\Large y'' = \frac{{y'\sqrt {{x^2} + 1} - y\frac{x}{{\sqrt {{x^2} + 1} }}}}{{{x^2} + 1}}\]

OpenStudy (trojanpoem):

i think we will multiply by x^2 +1

OpenStudy (michele_laino):

now I substitute my formula for y', so we get:

OpenStudy (michele_laino):

\[\Large \begin{gathered} y'' = \frac{{y'\sqrt {{x^2} + 1} - y\frac{x}{{\sqrt {{x^2} + 1} }}}}{{{x^2} + 1}} = \hfill \\ \hfill \\ = \frac{{\frac{{4y}}{{\sqrt {{x^2} + 1} }}\sqrt {{x^2} + 1} - y\frac{x}{{\sqrt {{x^2} + 1} }}}}{{{x^2} + 1}} \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

and after a simplification, I can write: \[\Large \begin{gathered} y'' = 4\frac{{y'\sqrt {{x^2} + 1} - y\frac{x}{{\sqrt {{x^2} + 1} }}}}{{{x^2} + 1}} = \hfill \\ \hfill \\ = 4\frac{{\frac{{4y}}{{\sqrt {{x^2} + 1} }}\sqrt {{x^2} + 1} - y\frac{x}{{\sqrt {{x^2} + 1} }}}}{{{x^2} + 1}} = \hfill \\ \hfill \\ = 4\frac{{4y\sqrt {{x^2} + 1} - yx}}{{{{\left( {{x^2} + 1} \right)}^{3/2}}}} \hfill \\ \end{gathered} \]

OpenStudy (trojanpoem):

How did you get the 4 ?

OpenStudy (michele_laino):

sorry I forgotten a factor before, namely 4, and the last steps are the right steps

OpenStudy (trojanpoem):

Oh, I see.

OpenStudy (michele_laino):

since the first derivative is: \[\frac{{4y}}{{\sqrt {{x^2} + 1} }}\]

OpenStudy (michele_laino):

now, we substitute our first and second derivatives into the left side of your equation

OpenStudy (michele_laino):

so we can write:

OpenStudy (trojanpoem):

By the way, can't we drive it into the proof directly ?

OpenStudy (michele_laino):

\[\Large \begin{gathered} \left( {{x^2} + 1} \right)y'' + xy' = \hfill \\ \hfill \\ = \left( {{x^2} + 1} \right)4\frac{{4y\sqrt {{x^2} + 1} - yx}}{{{{\left( {{x^2} + 1} \right)}^{3/2}}}} + x\frac{{4y}}{{\sqrt {{x^2} + 1} }} = \hfill \\ \hfill \\ = 4\frac{{4y\sqrt {{x^2} + 1} - yx}}{{\sqrt {{x^2} + 1} }} + \frac{{4xy}}{{\sqrt {{x^2} + 1} }} = \hfill \\ \hfill \\ = \frac{{16y\sqrt {{x^2} + 1} - 4xy + 4xy}}{{\sqrt {{x^2} + 1} }} = ...? \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

we have to substitute our y' and y'' into the left side of your equation, then we have to show that left side is equal to the rigt side

OpenStudy (trojanpoem):

I am doing it on paper now.

OpenStudy (trojanpoem):

I got (x^2 + 1)y'' = 16y - xy' / 4y

OpenStudy (michele_laino):

please complete my steps above

OpenStudy (trojanpoem):

Ok ,

OpenStudy (trojanpoem):

lol 16y

OpenStudy (michele_laino):

that's right!

OpenStudy (trojanpoem):

Thanks for the help , but still wanna inquiry why did I fail to get it be like the proof ?

OpenStudy (trojanpoem):

from 1st eqn sqrt(x^2+1) = 4y/ y'

OpenStudy (michele_laino):

in those type of differential equations, by my personal experience, I can say that at each step we have to rewrite a first derivative, or second derivative using the definition of the function, as we have made, otherwise the computations will become very difficult.

OpenStudy (trojanpoem):

Thanks ,I will do as you recommend. Can you help me with another one ? I will tag you in another question.

OpenStudy (michele_laino):

ok!